OpenStudy (anonymous):

A cup of coffee is made with boiling water at 100 C and stands in a room where the temperature is 20 C. The change in temperature, H in degrees C , with respect to time, t in minutes, is given by the following differential equation.

OpenStudy (anonymous):

the equation is \[\frac{ dH }{ dt }=-k(H-20)\]

OpenStudy (dumbcow):

separate variables \[\frac{dH}{H-20} = -k dt\] integrate both sides

OpenStudy (anonymous):

that's where i was having trouble. would the integral be ln(H-20)=-k^2/2?

OpenStudy (dumbcow):

ahh no right side you are integrating for variable "t" k is a constant --> ln(H-20) = -kt + C

OpenStudy (dumbcow):

always add the "+C" to right side

OpenStudy (anonymous):

oohhh okay. and the next part of the problem says "Solve this differential equation. If the coffee cools to 90 in 3 minutes, how long will it take to cool to 60 degrees?" would i be plugging something into the t and H to solve for C?

OpenStudy (dumbcow):

yes t = 3 , H = 90 you also know when t=0 H = 100

OpenStudy (dumbcow):

you have 2 constants to find ... both C and k C from initial conditions k from rate of cooling ... H = 90 after 3 min

OpenStudy (anonymous):

oh okay! i think i got what to do. thanks! :)

OpenStudy (dumbcow):

yw